The people of Vancouver cannot keep a secret. Upon being told a secret, a person from Vancouver will spend the next hour telling 3 friends the secret. In turn, those 3 friends will spend the next hour each telling 3 more people. This process continues and no one will tel someone who already knows the secret. If you tell a person from Vancouver a secret, how many people, excluding you, will know the secret in 7 hours?
Answer by amoresroy(361)

a = 3
r = 3
n = 7
Sum = a (1 - r^n)/((1-r
= 3 (1 - 3^7)/ (1-3)
= 3 (1-2187)/ (1-3)
= 3 (-2186)/(-2)
= 3,279
3,279 people will know the secret in 7 hours